OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a focus at (0, -10) and a directrix at y = 10

OpenStudy (anonymous):

@jim_thompson5910 HALP

jimthompson5910 (jim_thompson5910):

how far is the point (0, -10) away from the directrix y = 10

OpenStudy (anonymous):

20

jimthompson5910 (jim_thompson5910):

cut this in half to get 20/2 = 10

jimthompson5910 (jim_thompson5910):

the vertex will be 10 units away from the focus AND be 10 units away from the directrix

jimthompson5910 (jim_thompson5910):

so the vertex will lie on the midpoint of the segment connecting the focus and the directrix (this segment is also perpendicular to the directrix)

jimthompson5910 (jim_thompson5910):

so where is the vertex?

OpenStudy (anonymous):

0,0

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

the vertex in general is (h,k) so if the vertex is (0,0), then (h,k) = 0 h = 0 k = 0

jimthompson5910 (jim_thompson5910):

the distance from the vertex to the focus (call it p) is p = 10

jimthompson5910 (jim_thompson5910):

4p(y - k) = (x - h)^2 4*10(y - 0) = (x - 0)^2 ... plug in the three values just found 40y = x^2 y = (1/40)*x^2

jimthompson5910 (jim_thompson5910):

oh wait, i have my signs mixed up

OpenStudy (anonymous):

THANKS YOU'RE A GREAT MAN @jim_thompson5910

jimthompson5910 (jim_thompson5910):

it should be y = -(1/40)*x^2 because the parabola opens downward

OpenStudy (anonymous):

If the telescope is 1 m deep and 12 m wide, how far is the focus from the vertex? ONE MORE

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

let me think

jimthompson5910 (jim_thompson5910):

ok you'll have to do a drawing of a parabola

jimthompson5910 (jim_thompson5910):

but you'll have a parabola going through these 3 points (0,0) (6,1) (-6,1)

OpenStudy (anonymous):

done

jimthompson5910 (jim_thompson5910):

ok great

jimthompson5910 (jim_thompson5910):

The point (0,0) is the vertex. We'll use (6,1) as another point y = a(x-h)^2 + k y = a(x-0)^2 + 0 y = ax^2 1 = a(6)^2 1 = 36a a = 1/36 So the equation of the parabola is y = (1/36)x^2 ---------------------------------------------------- Now convert it to 4p(y - k) = (x - h)^2 form y = (1/36)x^2 y - 0 = (1/36)(x - 0)^2 36(y - 0) = (x - 0)^2 4*9(y - 0) = (x - 0)^2 It's now in 4p(y - k) = (x - h)^2 form where h = 0 k = 0 p = 9 So the distance from the vertex to the focus is 9 meters

OpenStudy (anonymous):

yay thanks so much!

jimthompson5910 (jim_thompson5910):

np